What mass of oxygen is needed to produce 75.0 grams of NO2?
1
Expert's answer
2019-09-04T04:46:34-0400

2 atoms of oxygen are required for producing 1 "NO_2" molecule

Molar mass of "NO_2=14+(2\times16)=46\frac{g}{mol}"

So,

"46 g" "NO_2" is produced by "32g" oxygen

"75g" "NO_2" is produced by "\frac{32}{46}\times 75=52.174g" oxygen

mass of oxygen needed to produce 75.0 grams of "NO_2=52.174g"